### v1

# Teaching Quantum Computing through Lab-Integrated Learning: Bridging Conceptual and Computational Understanding

Abstract.
Quantum computing education requires students to move beyond classical programming intuitions related to state, determinism, and debugging, and to develop reasoning skills grounded in probability, measurement, and interference. This paper reports on the design and delivery of a combined undergraduate and graduate course at Louisiana State University that employed a lab-integrated learning model to support conceptual change and progressive understanding. The course paired lectures with weekly programming labs that served as environments for experimentation and reflection. These labs enabled students to confront misconceptions and refine their mental models through direct observation and evidence-based reasoning. Instruction began with Quantum Without Linear Algebra (QWLA), which introduced core concepts such as superposition and entanglement through intuitive, dictionary representations. The course then transitioned to IBM Qiskit, which provided a professional framework for circuit design, noise simulation, and algorithm implementation. Analysis of student work and feedback indicated that hands-on experimentation improved confidence, conceptual clarity, and fluency across representations. At the same time, it revealed persistent challenges in debugging, reasoning about measurement, and understanding probabilistic outcomes. This paper presents the course structure, instructional strategies, and lessons learned, and argues that lab-integrated learning offers an effective and accessible approach to teaching quantum computing in computer science education.
quantum computing, quantum programming, curriculum, case study
journalyear: 2025copyright: none
1. Introduction
Quantum computing (QC) is redefining the limits of computation by applying quantum-mechanical principles such as superposition, entanglement, and interference to process information in fundamentally new ways. These principles enable algorithms that can outperform classical methods on problems such as integer factorization (Shor, 1994), unstructured search (Grover, 1996), and molecular simulation (Low et al., 2019). As both research and industry accelerate progress toward scalable quantum hardware, the demand for a workforce proficient in both the theory and practice of quantum computing is growing rapidly.
Meeting this demand presents a distinct challenge for computer science education. Most existing quantum computing courses are offered at the graduate level and focus heavily on mathematical derivations or the physical implementation of qubits. While these topics provide theoretical depth, they often limit accessibility for students with backgrounds in programming and algorithmic reasoning. On the other hand, courses that focus solely on software may encourage superficial reproduction of circuits without fostering conceptual understanding. Addressing this gap requires pedagogical approaches that integrate experimentation with conceptual learning and link abstract ideas to computational representations.
At Louisiana State University (lsu, 025c), we designed and taught a course titled Introduction to Quantum Computing, open to both undergraduate and graduate computer science students. The course follows a lab-integrated learning model that combines conceptual instruction with hands-on programming every week. Students begin with Quantum Without Linear Algebra (QWLA) (Albarghouthi, 2025), a lightweight simulation framework that supports intuitive engagement with core quantum ideas without requiring prior knowledge of matrix algebra. Once students build foundational fluency, the course transitions to IBM Qiskit (Wille et al., 2019) for constructing and analyzing quantum circuits and algorithms. Each weekly lab targets a specific learning objective, such as superposition, entanglement, or noise mitigation, and reinforces it through executable experiments and guided reflection.
This instructional approach aims to balance accessibility with intellectual rigor. By grounding theoretical concepts in programming and experimentation, students develop an understanding of quantum program behavior, error sources, and the relationship between abstractions and hardware execution. In this report, we share our experience designing and teaching this lab-integrated course. We describe the motivations behind its instructional design, present the core components of the curriculum, and reflect on student outcomes and feedback. Our goal is to offer a replicable framework that integrates conceptual learning and programming practice in quantum computing education.
2. Background
Traditional approaches to teaching quantum computing often emphasize theoretical foundations, typically delivered through physics or mathematics departments. Courses offered in physics programs tend to focus on the behavior of quantum systems and hardware implementation (cal, 2019; ber, 2014), while those in mathematics departments emphasize formal aspects of information theory and linear algebra (ber, 2007; wat, 2019). In both cases, assessment usually relies on written assignments, exams, and final projects presented as reports or oral presentations on selected algorithms.
These courses are generally targeted at graduate students. Undergraduate quantum computing offerings within computer science departments remain relatively uncommon (ut-, 2017; ucl, 2019). However, there has been increasing interest in developing programming-oriented and software-focused quantum computing courses that prioritize algorithmic thinking, practical experimentation, and computational abstraction. Notable examples include the University of Washington’s course Introduction to Quantum Computing and Quantum Programming in Q# (uw, 2019; Mykhailova and Svore, 2020), and the work by Seegerer and others (Seegerer et al., 2021), which explores how core ideas such as superposition, entanglement, and quantum algorithms can be taught through computing education frameworks.
These efforts represent a meaningful shift toward integrating theory with practice. Programming-based instruction has shown that active experimentation can make quantum concepts more accessible to computer science students. Nevertheless, many current courses still lack dedicated laboratory components that allow students to formulate hypotheses, visualize results, and debug their understanding through code. By comparison, classical computing education often begins with hands-on programming, then progresses toward algorithms and theory through iterative practice. Quantum computing education stands to benefit from a similar approach, where theoretical concepts are grounded in coding and experimentation rather than introduced as purely abstract content.
3. Theoretical Foundation
The course design was guided by three perspectives: (1) the core ideas that define conceptual understanding in quantum computing, (2) a learning process centered on conceptual change through experimentation and reflection, and (3) a progressive framework that moves from QWLA to Qiskit, linking intuitive understanding with practical programming experience.
3.1. Core Ideas in QC Education
Seegerer and others (Seegerer et al., 2021) identify five core ideas that underpin conceptual understanding in quantum computing: superposition, entanglement, quantum computation, quantum algorithms, and quantum cryptography. These ideas define the mental models students must develop to think quantum mechanically. For example, superposition introduces probabilistic reasoning about state, entanglement extends beyond classical correlation, and quantum algorithms demonstrate how interference and measurement can be used to amplify correct outcomes.
Traditional instruction often presents these ideas through dense mathematics or physical analogies, which can obscure their computational relevance. Our course builds on this conceptual framework by embedding each idea within a lab sequence that combines explanation, simulation, and hands-on experimentation. Students first engage with these concepts intuitively, then explore their formal structure through programming exercises that reveal their computational meaning.
3.2. Conceptual Change and Evolution
Learning quantum computing involves more than absorbing new content; it requires reconfiguring existing knowledge. Students often bring classical intuitions about computation and probability that conflict with quantum principles. For instance, they may believe that measurement reveals a pre-existing value rather than collapsing a probabilistic state, or assume that correlations imply hidden communication instead of entanglement.
Figure 1. Comparison of classical and quantum reasoning. In the classical view, a bit is either 0 or 1, while in the quantum view, a qubit can exist in a superposition of both states until measurement. The figure illustrates the conceptual shift students experience when learning to reason about state, probability, and measurement in quantum programs.
Figure 1 illustrates this shift in perspective by comparing classical and quantum models of computation. In classical reasoning, a bit exists in a definite state, either 0 or 1. A qubit can exist in a superposition of both states until it is measured. Measurement collapses the superposition, producing a single outcome. The figure highlights how classical views based on determinism and discrete probability differ from quantum reasoning, where computation depends on amplitudes, interference, and probabilistic results.
The course supports this conceptual evolution by placing students in situations where classical assumptions are challenged through direct experimentation. Each lab provides opportunities for students to observe mismatches between their expectations and actual quantum behavior, which encourages reflection and refinement of their mental models. Experiences such as seeing measurement collapse or encountering non-classical correlations in code serve as turning points that deepen understanding.
This process reflects the broader concept of conceptual change in science education. Students move from initial, often classical, interpretations to more accurate quantum models. In this course, laboratory-based exploration functions as the primary mechanism for facilitating that transition, linking abstract theoretical ideas to observable computational outcomes.
3.3. Progressive Shift to Quantum Programming
Teaching quantum programming presents several fundamental challenges that differ from classical programming. (1) Quantum basis states cannot be directly inspected or debugged, because measurement collapses the state and eliminates traditional feedback mechanisms. (2) Students must reason across multiple representations, including Dirac notation, matrices, Bloch sphere visualizations, and circuit diagrams, with each offering a distinct perspective on quantum operations. (3) Entanglement introduces non-local dependencies, which undermine the modular reasoning that is central to classical programming. (4) Correctness cannot be inferred from single runs. Due to probabilistic outputs and noise, repeated trials and comparison with theoretical expectations are required.
Because of these challenges, the course uses a slow and guided transition between conceptual and programming frameworks. The course begins with QWLA, which offers an accessible introduction for students with limited backgrounds in quantum mechanics. QWLA enables students to model key phenomena, such as superposition, measurement, and entanglement, using simple Python constructs. This stage builds foundational intuition without requiring advanced mathematics, allowing all students to meaningfully engage with early course content.
Following the introductory phase, the course transitions to IBM Qiskit, which serves as the primary platform for quantum algorithm design, circuit implementation, and experimental analysis. Qiskit offers a professional-grade development environment that enables students to translate conceptual understanding into executable quantum programs. Through systematic comparison of simulation outputs with theoretical predictions, students develop the ability to interpret algorithmic behavior, assess the influence of noise, and reason about computational performance.
This progression from intuitive modeling in QWLA to formal development in Qiskit demonstrates how effective scaffolding can lower entry barriers while guiding students toward conceptual maturity. The integration of core ideas, conceptual evolution, and graduated tooling forms the theoretical foundation of our lab-integrated approach to quantum computing education.
4. Instructional Design
Building on the theoretical foundations described earlier, the instructional design emphasized accessibility, experimentation, and progressive abstraction. The goal was not only to introduce the mechanics of quantum programming but also to help students internalize the reasoning behind quantum phenomena through guided observation and reflection.
4.1. Learning Approach
The course followed a lab-integrated learning model in which lectures and labs were closely connected. Each week included one 80-minute lecture and one 80-minute lab session. Lectures introduced key ideas and demonstrated important algorithms and principles. Labs provided guided opportunities for students to apply, test, and extend those ideas. This structure enabled conceptual understanding and computational skills to develop in parallel.
To make the course accessible, instruction began with QWLA, which allowed students to explore foundational concepts such as superposition, measurement, and entanglement using intuitive, code-based representations. Initially, quantum states were introduced as Python dictionaries that mapped basis states to amplitudes, providing a simple and concrete way to reason about superposition and measurement without requiring matrix algebra. This representation helped students visualize state transformations and probabilities in a familiar programming context. After students became comfortable with these behaviors, the course transitioned to IBM Qiskit, which served as the main platform for designing algorithms, building circuits, and running experiments. This progression allowed students to move from informal reasoning to formal implementation, connecting conceptual learning with practical skill development.
Each lab was explicitly aligned with the concepts introduced in the corresponding lecture. For example, after covering entanglement and measurement correlations, students built and tested Bell-state circuits, analyzed measurement results, and explored the effects of noise and decoherence. These lab activities served as formative checkpoints that helped students connect code and circuits with quantum behavior.
4.2. Course Structure
The course was offered over sixteen weeks to both undergraduate and graduate students who had completed an introductory algorithms course. No prior background in quantum mechanics was required. The curriculum (Table 1) advanced gradually from foundational concepts to more complex algorithms and applications, balancing theoretical instruction with experimental practice.
Table 1. Alignment of Lecture Topics and Lab Activities
Lecture Topics Laboratory Focus
Qubits, superposition, and measurement Exploring single-qubit states using QWLA
Entanglement and correlations Building and testing Bell states
Phase and interference Phase kickback
Quantum circuits and the Deutsch-Jozsa algorithm IBM Qiskit circuits and implementing Deutsch-Jozsa
Grover’s search Comparing quantum and classical search performance
Noise, and error mitigation Simulating noise models in Qiskit
Quantum Fourier Transform and Phase Estimation Coding QFT and analyzing frequency outputs
Shor’s algorithm and applications Factorization and scaling analysis
This structure reinforced the idea that each theoretical topic should have a corresponding experimental activity. The gradual increase in complexity helped students build algorithmic thinking alongside scientific reasoning about quantum systems.
4.3. Assessment and Learning Outcomes
Assessment combined formative and summative methods to evaluate both conceptual understanding and technical proficiency. Laboratory assignments were assessed automatically using provided testing scripts. The final project required teams of two or three students to design, implement, and evaluate a quantum algorithm or application. Project presentations were evaluated manually based on implementation quality, conceptual insight, and clarity of communication. Example projects included executing Grover’s search under noisy conditions and developing small-scale quantum optimization tasks. Additional assessment components included quizzes, a midterm examination, and class participation.
By the end of the course, students were expected to:
  • Explain core quantum concepts such as superposition, entanglement, and measurement using both descriptive and computational reasoning.
  • Design and implement quantum circuits and algorithms using Qiskit.
  • Analyze the effects of noise, error, and decoherence on quantum computation.
  • Relate theoretical operations to their practical implementations and observable outcomes.
The alignment between lectures and labs, along with the gradual progression from intuitive simulation to professional development tools, created a learning environment in which students engaged with quantum computing through experimentation, analysis, and reflection.
5. Implementation and Delivery
The course was delivered over a sixteen-week semester at LSU in Fall 2025. It enrolled 9 undergraduate and 6 graduate students in computer science and software engineering. The class met twice weekly, with one 80-minute lecture focused on conceptual foundations and algorithm design, followed by an 80-minute lab session dedicated to hands-on programming and experimentation.
5.1. Lecture Format
Lectures introduced the theoretical foundations and computational interpretations of each topic. Visualizations and live coding demonstrations were used extensively to connect abstract concepts with executable code. For example, superposition and interference were first illustrated using diagrams and probability plots, followed by step-by-step implementation in live code.
All instructional materials, including lecture slides, annotated code, and Jupyter notebooks, were made available on the course website after each session. This allowed students to review content at their own pace. The lecture format emphasized intuitive understanding first, then practical implementation, and formal mathematics when necessary. Students were encouraged to discuss interpretations of quantum behavior before encountering formal representations in code.
5.2. Laboratory Sessions
Following the Progressive transition to quantum programming approach (Sec 3.3), the lab sessions gradually increase in complexity as students transition from intuitive experimentation to full algorithmic implementation. Each lab extends lecture content and serves as the primary space for hands-on learning. This structure allows students to move seamlessly between reasoning about quantum phenomena and validating their understanding through simulation.
In the early weeks, labs used QWLA to introduce fundamental concepts such as qubits, superposition, and measurement. These concepts were explored through dictionary-based code representations, which allowed students to visualize state transformations without relying on matrix algebra. This phase helped students develop early intuition and confidence in quantum programming. Starting in week four, IBM Qiskit became the primary development platform. Students constructed circuits, simulated algorithms, and analyzed measurement outcomes. Students were encouraged to use an automated testing framework during labs for immediate feedback on correctness.
5.3. Projects
The course concluded with a team-based final project. Students worked in pairs to investigate a topic related to quantum algorithms or applications. Example topics included comparing classical and quantum search methods, analyzing noise and error mitigation strategies, and implementing small-scale versions of the Quantum Approximate Optimization Algorithm (QAOA).
Each team maintained a GitHub repository containing their code, documentation, and analysis. Projects were evaluated based on implementation correctness, clarity of communication, originality, and depth of reflection. The final two class sessions were devoted to student presentations, where each group demonstrated their project and discussed the challenges encountered during development.
5.4. Student Interaction and Support
Collaboration and discussion were integral to the course. Students were encouraged to work together during labs, share debugging strategies, and explain ideas to one another. Office hours and an online discussion forum were available for additional support. The teaching assistant provided in-lab guidance and helped students interpret simulation outputs.
Several students, especially those without prior exposure to quantum mechanics, initially expressed concern about their readiness. However, the iterative structure of the labs and the emphasis on experimentation helped reduce this anxiety. By the fourth week of the semester, most students reported increased confidence in reasoning about quantum behavior and verifying their understanding through code.
6. Lessons Learned
The first offering of this course provided valuable insights into how students engage with quantum computing when instruction emphasizes iterative experimentation and conceptual reflection. Feedback was collected through end-of-semester surveys, written reflections, informal interviews, and observations by the instructional staff.
Lowering Barriers to Entry. The early use of QWLA significantly lowered the barrier to entry for students with no prior exposure to quantum mechanics. By introducing qubits, superposition, and measurement through intuitive, dictionary-based code, students were able to explore fundamental ideas without requiring linear algebra or complex syntax. This helped demystify quantum phenomena and allowed students to focus on conceptual understanding before engaging with formal representations.
The subsequent transition to IBM Qiskit introduced a professional-grade platform that brought students closer to real-world quantum programming. Many students reported that encountering the same concepts in Qiskit after working with them in QWLA reinforced their understanding and illustrated how conceptual models map onto executable circuits.
Classical to. Quantum Thinking: Concepts and Programming. One of the most persistent challenges students faced was adjusting their mental model from classical to quantum computation. Many approached quantum circuits with expectations shaped by classical programming, such as the ability to inspect intermediate states or trace execution step-by-step. These assumptions led to confusion around key concepts like state collapse, probabilistic execution, and interference.
To address these misconceptions, several labs included direct comparisons between classical and quantum approaches. For example, search and parity-checking tasks were implemented in both paradigms to highlight how probability and amplitude interference replace deterministic control flow. These experiences helped students reframe their understanding of computation and appreciate the unique constraints and affordances of the quantum model.
Building Confidence Through Experimentation. Weekly labs played a central role in building student confidence. The ability to run small programs and observe immediate results allowed students to test their understanding incrementally. Exercises that produced clear visual or numerical feedback, such as state vector plots or Bell-state measurements, were especially effective in reinforcing key ideas. Students noted that these experiments made abstract concepts more concrete and improved their ability to explain quantum behavior in their own words.
This feedback loop between intuition and execution was critical in supporting students who were initially unsure about their ability to grasp quantum mechanics. The hands-on approach allowed students to shift from passive observation to active discovery.
Collaboration and Peer Learning. Collaborative learning was a consistent theme throughout the course. Students frequently worked in pairs during labs, exchanged debugging strategies, and discussed quantum behavior informally. Peer interaction played a key role in surfacing and resolving misconceptions, especially during live coding sessions.
The final projects extended this collaboration. Students selected open-ended topics such as quantum optimization or machine learning applications, which encouraged them to synthesize ideas from across the course. Teams documented their work in public GitHub repositories and presented their findings in the final sessions. These projects demonstrated both conceptual mastery and creative adaptation beyond the core syllabus.
Managing Complexity and Workload. While most students found the lab submissions rewarding, several noted that the workload increased significantly during the second half of the semester. Labs involving Grover’s algorithm, noise modeling, and Shor’s algorithm required more time and debugging effort than anticipated.
To reduce cognitive load in future iterations, later labs will include more structured scaffolding, such as pre-written starter code and step-by-step validation checkpoints. Students also suggested brief pre-lab tutorials on Qiskit and debugging strategies. These adjustments are intended to preserve the rigor of the lab content while improving the overall balance of effort across the semester.
Instructional Insights. From an instructional perspective, the lab-integrated design increased preparation time but significantly enhanced student engagement. The continuous alignment of theory with hands-on practice created opportunities to address misconceptions as they emerged. Common points of confusion included the probabilistic nature of measurement, the distinction between amplitudes and probabilities, and reasoning about phase.
The iterative lab structure proved especially effective for enabling early intervention and reinforcing concepts before advanced topics were introduced. The course experience reaffirmed the importance of coupling conceptual explanation with experimentation in helping students reason meaningfully about quantum behavior.
7. Conclusions and Future Work
The initial offering of this course demonstrated that lab-integrated instruction can make quantum computing both accessible and intellectually engaging for computer science students. Pairing concise lectures with weekly programming labs enabled students to connect abstract concepts such as superposition, entanglement, and interference with observable behavior in code. The structured progression from QWLA to Qiskit supported the development of conceptual understanding and increased confidence in using professional tools.
Students consistently reported that hands-on experimentation clarified difficult concepts and improved their ability to explain quantum behavior. At the same time, the course revealed specific areas where students encountered difficulties, particularly in debugging probabilistic systems, interpreting measurement outcomes, and adjusting to the absence of intermediate state visibility. These challenges highlighted the cognitive transition required when moving from classical to quantum programming.
To address these issues, future offerings of the course will include refinements to lab materials to better balance complexity and workload. Planned improvements include short pre-lab tutorials on syntax and debugging strategies, along with enhanced feedback and visualization tools.